# PORTULAN ExtraGLUE Datasets and Models: Kick-starting a Benchmark for the Neural Processing of Portuguese

### v2

@@ 3.3 Inference tasks
GLUE and SuperGLUE also include an expert-constructed diagnostic dataset covering different types of linguistic phenomena. Broadcoverage Diagnostics (AX)GS (Wang et al.,, 2018) is a Natural Language Inference (NLI) task designed to test models across a wide spectrum of linguistic, commonsense, and world knowledge; each instance contains a sentence pair labeled with entailment or not entailment. Winogender Schema Diagnostics (AX)S (Rudinger et al.,, 2018) is a similar task, designed to measure gender bias, where each premise sentence includes a male or female pronoun and a hypothesis includes a possible referent for the pronoun.
@@ 4.2 Gendered nouns and pronoun resolution
English common nouns do not express grammatical gender. On the other hand, Portuguese common nouns do and are used with corresponding gendered determiners (as opposed to English gender-neutral the or a). This exacerbates the difficulty of properly addressing pronoun reference resolution, given that third-person singular pronouns (and also plural in Portuguese) are gendered. Tasks specifically dealing with pronoun resolution or evaluating the gender robustness of language models are thus prone to corruption via MT. These include WNLI, WSC, and AX. While we provide translated versions of WNLI and AX, we conduct error analysis to diagnose the quality level of their Portuguese versions.
@@
AX focuses on gender bias, explicitly combining both concerns expressed above. For instance, the hypothesis “The investigator tried to get in contact” is translated into Portuguese as “O investigador tentou entrar em contacto”; its possibly accompanying premises “The investigator wanted to interview the witness in person, but [he | she] was unable to get in contact” are translated into Portuguese as “O investigador queria entrevistar a testemunha pessoalmente, mas não conseguiu entrar em contacto com ela” (for he), or to “O investigador queria entrevistar a testemunha pessoalmente, mas ela não conseguiu entrar em contacto” (for she). In the latter case, limiting the possible referents of pronoun ela (she) – the only feminine noun is testemunha (witness), since investigador (investigator) is masculine in Portuguese – renders the label wrong, as it should be changed to .
@@ 4.4 Machine-translated tasks
The set of datasets that have been translated and are part of PORTULAN ExtraGLUE 555Made available at [url temporarily removed to preserve anonymity during review]. are included in Table 1. As mentioned in Sections 4.1 and 4.2, we leave out the CoLA and WSC datasets.
| Task | Train | Val | Test | Tokens (en) | Version | Tokens (pt) | mt | lab | low |
@@
The main translation problems we have observed concern pronoun resolution or gender issues (as already emphasized in Section 4.2), idiomatic expressions, inconsistent translations in pairs of sentences, and a few cases of ‘hallucinations,’ among other problematic mistranslations. In some cases, these problems have an impact on the correctness of the labels (mainly in WNLI and AX); in other cases, they mostly imply a dataset of lesser quality (such as in AX and BoolQ). In the specific case of AX, even when the translation is correct, it does not do justice to the nature of the task, which loses its purpose (e.g., his/her translate the same way to Portuguese).
@@ Acknowledgements
This research was partially supported by: PORTULAN CLARIN — Research Infrastructure for the Science and Technology of Language, funded by Lisboa 2020, Alentejo 2020 and FCT (PINFRA/22117/2016); ACCELERAT.AI - Multilingual Intelligent Contact Centers, funded by IAPMEI (C625734525-00462629); ALBERTINA - Foundation Encoder Model for Portuguese and AI, funded by FCT (CPCA-IAC/AV/478394/2022); and Base Funding (UIDB/00027/2020) and Programmatic Funding (UIDP/00027/2020) of the Artificial Intelligence and Computer Science Laboratory (LIACC) funded by national funds through FCT/MCTES (PIDDAC); Bernardo Leite is supported by a PhD studentship (with reference 2021.05432.BD), funded by Fundação para a Ciência e a Tecnologia (FCT).

### v3

@@ 3.3 Inference tasks
GLUE and SuperGLUE also include an expert-constructed diagnostic dataset covering different types of linguistic phenomena. Broadcoverage Diagnostics (AXb)GS (Wang et al.,, 2018) is a Natural Language Inference (NLI) task designed to test models across a wide spectrum of linguistic, commonsense, and world knowledge; each instance contains a sentence pair labeled with entailment or not entailment. Winogender Schema Diagnostics (AXg)S (Rudinger et al.,, 2018) is a similar task, designed to measure gender bias, where each premise sentence includes a male or female pronoun and a hypothesis includes a possible referent for the pronoun.
@@ 4.2 Gendered nouns and pronoun resolution
English common nouns do not express grammatical gender. On the other hand, Portuguese common nouns do and are used with corresponding gendered determiners (as opposed to English gender-neutral the or a). This exacerbates the difficulty of properly addressing pronoun reference resolution, given that third-person singular pronouns (and also plural in Portuguese) are gendered. Tasks specifically dealing with pronoun resolution or evaluating the gender robustness of language models are thus prone to corruption via MT. These include WNLI, WSC, and AXg. While we provide translated versions of WNLI and AXg, we conduct error analysis to diagnose the quality level of their Portuguese versions.
@@
AXg focuses on gender bias, explicitly combining both concerns expressed above. For instance, the hypothesis “The investigator tried to get in contact” is translated into Portuguese as “O investigador tentou entrar em contacto”; its possibly accompanying premises “The investigator wanted to interview the witness in person, but [he | she] was unable to get in contact” are translated into Portuguese as “O investigador queria entrevistar a testemunha pessoalmente, mas não conseguiu entrar em contacto com ela” (for he), or to “O investigador queria entrevistar a testemunha pessoalmente, mas ela não conseguiu entrar em contacto” (for she). In the latter case, limiting the possible referents of pronoun ela (she) – the only feminine noun is testemunha (witness), since investigador (investigator) is masculine in Portuguese – renders the label wrong, as it should be changed to .
@@ 4.4 Machine-translated tasks
The set of datasets that have been translated and are part of PORTULAN ExtraGLUE 555Made available at https://huggingface.co/datasets/PORTULAN/extraglue. are included in Table 1. As mentioned in Sections 4.1 and 4.2, we leave out the CoLA and WSC datasets.
| Task | Train | Val | Test | Tokens (en) | Version | Tokens (pt) | mte | labe | lowq |
@@
The main translation problems we have observed concern pronoun resolution or gender issues (as already emphasized in Section 4.2), idiomatic expressions, inconsistent translations in pairs of sentences, and a few cases of ‘hallucinations,’ among other problematic mistranslations. In some cases, these problems have an impact on the correctness of the labels (mainly in WNLI and AXg); in other cases, they mostly imply a dataset of lesser quality (such as in AXb and BoolQ). In the specific case of AXg, even when the translation is correct, it does not do justice to the nature of the task, which loses its purpose (e.g., his/her translate the same way to Portuguese).
@@ Acknowledgements
This research was partially supported by: PORTULAN CLARIN — Research Infrastructure for the Science and Technology of Language, funded by Lisboa 2020, Alentejo 2020 and FCT (PINFRA/22117/2016); ACCELERAT.AI - Multilingual Intelligent Contact Centers, funded by IAPMEI (C625734525-00462629); ALBERTINA - Foundation Encoder Model for Portuguese and AI, funded by FCT (CPCA-IAC/AV/478394/2022); Base Funding (UIDB/00027/2020) and Programmatic Funding (UIDP/00027/2020) of the Artificial Intelligence and Computer Science Laboratory (LIACC) funded by national funds through FCT/MCTES (PIDDAC); Bernardo Leite is supported by a PhD studentship (with reference 2021.05432.BD), funded by Fundação para a Ciência e a Tecnologia (FCT).